OpenStudy (kaliyah0524):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (kaliyah0524):

Im not asking for the answer. I just need help working it out

OpenStudy (johnweldon1993):

\[\large d = ut + 0.5at^2\] Well the first thing you want to do, is isolate the terms that are being multiplied to the 'a' so we need to subtract \(\large ut\) from both sides of this equation \[\large d - ut = 0.5at^2\] What would be next? (keep in mind, the only thing being squared here, is the 't')

OpenStudy (kaliyah0524):

To subtract the 0.5?

OpenStudy (johnweldon1993):

Well no, since 0.5 is being multiplied to the 'a' we need to cancel that out by DIVIDING by 0.5 So we actually divide everything by 0.5 \[\large \frac{d - ut}{0.5} = at^2\] make sense? and so what would we do next?

OpenStudy (kaliyah0524):

Not sure

OpenStudy (johnweldon1993):

Well it might be easier to write this as \[\large \frac{d - ut}{0.5} = a \times t^2\] Now we can clearly see that t^2 is being multiplied to 'a'...and so we do the sae thing we did with the 0.5...we divide it \[\large \frac{\frac{d - ut}{0.5}}{\frac{t^2}{1}} = \frac{d - ut}{0.5t^2}\]

OpenStudy (kaliyah0524):

Confused :/

OpenStudy (kaliyah0524):

I divide t^2 but D-ut/0.5 ?

OpenStudy (kaliyah0524):

by*

OpenStudy (kaliyah0524):

So the answer would be A=D-UT/0.5/t^2 ?

OpenStudy (kaliyah0524):

@johnweldon1993